# CTISum: A New Benchmark Dataset For Cyber Threat Intelligence Summarization

### v1

@@ Abstract.
Cyber Threat Intelligence (CTI) summarization task requires the system to generate concise and accurate highlights from raw intelligence data, which plays an important role in providing decision-makers with crucial information to quickly detect and respond to cyber threats in the cybersecurity domain. However, efficient techniques for summarizing CTI reports, including facts, analytical insights, attack processes, etc., have largely been unexplored, primarily due to the lack of available dataset. To this end, we present CTISum, a new benchmark for CTI summarization task. Considering the importance of attack process, a novel fine-grained subtask of attack process summarization is proposed to enable defenders to assess risk, identify security gaps, vulnerabilities, and so on. Specifically, we first design a multi-stage annotation pipeline to gather and annotate the CTI data, and then benchmark the CTISum with a collection of extractive and abstractive summarization methods. Experimental results show that current state-of-the-art models exhibit limitations when applied to CTISum, underscoring the fact that automatically producing concise summaries of CTI reports remains an open research challenge.
Intelligence analysis, Summarization, Benchmark and dataset
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Summarizationccs: Computing methodologies Natural language generation
1. Introduction
@@
Cyber Threat Intelligence (CTI), also known as threat intelligence, is knowledge, skills and experience-based information concerning the occurrence and assessment of both cyber and physical threats as well as threat actors (Jo et al., 2022; of England, 2016). The CTI data is naturally unstructured, voluminous, fragmented and technical, making it difficult for analysts to efficiently locate the most relevant and high-value intelligence. Therefore, it becomes crucial to automatically summarize the knowledge contained in CTI reports, which could help analysts simply identify events, patterns, cyber attacks and conclusions. Furthermore, CTI summarization has broad applicability across domains like cybersecurity (Schatz et al., 2017), military intelligence (van Tilborg, 2005), technical alerts, threat modeling, and more, where complicated information needs to be distilled into concise insights.
@@
Figure 1. An example in the CTISum. Blue arrow and orange arrow indicate the CTIS and APS tasks, respectively. Red font represents the key words.
@@
With the ongoing advancement of summarization techniques, numerous datasets have emerged, such as CNN/DailyMail (Nallapati et al., 2016) and XSum (Narayan et al., 2018), among others. In addition, the introduction of these datasets further drives the development of associated technologies like PGNet (See et al., 2017), BART (Lewis et al., 2020), BRIO (Liu et al., 2022), etc. Furthermore, domain-specific summarization researches have gradually arisen. Examples include biomedicine (Zhu et al., 2023), finance (Mukherjee et al., 2022), law (Shukla et al., 2022), etc. This further facilitates various applications involving aiding medical decision-making, generating financial reports, and summarizing legal documents. The boom of domain-specific summarization promotes the understanding of domain terminology and context, which is crucial for generating high-quality summaries for further research.
@@
Despite remarkable progress made in automatic text summarization, effective methods for summarizing CTI reports in the cybersecurity domain remain largely unexplored. This is primarily due to the lack of available dataset. The unique characteristics of CTI reports, such as technical jargon, evolving threat landscapes, longer reports and fragmented data information, make it challenging to create an annotation, which in turn hinders the development of robust models and evaluation benchmarks. However, recent advances in Large Language Models (LLMs) (Hsiao and Collins, 2023; Cui et al., 2024; OpenAI, 2023; Touvron et al., 2023; Claude, 2023b, a) have shown promising text understanding and generation capabilities. Therefore, this raises the question: How might we construct a high-quality CTI summarization dataset with the assistance of LLMs?
@@
In this paper, we construct a new benchmark CTISum based on threat intelligence obtained from different sources in the cybersecurity domain. In addition to the CTI Summarization (CTIS) task, considering the importance of attack process, we propose a novel fine-grained subtask, Attack Process Summarization (APS), to enable defenders to quickly understand reported attack behaviors, assess risks and identify security vulnerabilities. Specifically, a multi-stage annotation pipeline is designed to preprocess CTI reports and obtain annotation data with the assistance of LLMs, which includes the data collection stage, parsing & cleaning stage, prompt schema stage and intelligence summarization stage. To further keep the high quality of the CTISum, we additionally employ expert reviews for double checking (details can be seen in Section 3.1). With proper data, automated systems can help synthesize lengthy CTI reports into concise and accurate summaries for different intelligence analysts. A sampled report and reference summary in CTISum are shown in Figure 1. The system not only requires to make a general summary of the full document (task 1), but also needs the capability to focus on and generate the attack process (task 2). After cleaning and human checking, CTISum obtains 1,345 documents and corresponding summaries.
@@
What makes CTISum a challenging dataset can be presented as follows. First, the average document length is about 2,865 words. Traditional deep learning techniques are incapable of processing documents surpassing 512/1024 tokens in length (after tokenization), such as GPT1(Radford et al., 2018), T5(Raffel et al., 2020), BART(Lewis et al., 2020). Therefore, although achieving state-of-the-art (SOTA) performances on short-document summarization datasets, these models cannot be effectively transferred to CTISum. LLMs can deal with longer input, however, fine-tuning the LLMs is resource and time consuming. Second, the document-to-summary compression ratio on the two tasks are 14.32 (2865.60 / 200.04) and 22.23 (2865.60/118.27), respectively (details in Table 1). The high compression ratio shows competitiveness with the current long document summarization datasets, requiring systems to be extremely precise in capturing only the most relevant facts from lengthy documents in a minimal number of words. Third, CTISum has another subtask, APS, which means the system should be capable of the ability to capture the fine-grained attack process described in CTI reports.
@@
Figure 2. The overview of the proposed multi-stage annotation pipeline for CTIS and APS tasks.
2. Related Work
2.1. Text Summarization
A considerable amount of existing researches in automatic text summarization (Rush et al., 2015; Nallapati et al., 2016, 2017; Celikyilmaz et al., 2018; Lebanoff et al., 2019; Zhong et al., 2020) have explored to improve summary generation for news articles, using popular datasets such as CNN/DailyMail (Nallapati et al., 2016), XSum (Narayan et al., 2018), etc. In addition, scientific document summarization has emerged as another critical area, with datasets based on ArXiv/PubMed (Cohan et al., 2018), and other academic sources. The main approaches cover two aspects. First, extractive methods (Erkan and Radev, 2004; Cheng and Lapata, 2016; Zhou et al., 2018; Liu, 2019; Liu and Lapata, 2019; Zhong et al., 2020) involve identifying and extracting key sentences or passages from the original text to form the summary. For instance, Liu and Lapata (2019) introduce a document-level encoder based on BERT and propose a general framework for summarization. Second, abstractive techniques focus on generating new sentences to capture the meaning of the source text. For example, sequence-to-sequence models with attention (Luong et al., 2015; Nallapati et al., 2017; Peng et al., 2021) are usually utilized to generate the probability distributions on the vocabulary. In addition, See et al. (2017) present a hybrid pointer generator architecture for abstractive summarization. And more recently, transformer-based models like T5 (Raffel et al., 2020), BART (Lewis et al., 2020) and Longformer(Beltagy et al., 2020) have shown advanced performance on the summarization task.
@@ 2.1. Text Summarization
2.2. Domain-specific Summarization
For the domain-specific summarization, there are many potential applications across different domains like government (Huang et al., 2021), finance (Mukherjee et al., 2022), court rulings (Shukla et al., 2022) and biomedicine (Zhu et al., 2023). For example, in the government domain, summarization systems can help process large volumes of regulatory text and legal documents. Huang et al. (2021) propose a novel efficient encoder-decoder attention with head-wise positional strides to effectively double the processed input sequence size. For the finance industry (Mukherjee et al., 2022), summarization of earnings reports, financial statements, and news can assist investors and analysts in making decisions. In addition, summarizing court rulings, case law, and litigation documents (Shukla et al., 2022) helps legal professionals in their work. Within biomedicine (Zhu et al., 2023), healthcare professionals can make informed decisions by using summaries of clinical trial findings, medical literature, and patient records. Different from the above domains, we make the first attempt to build the CTI summarization dataset in the cybersecurity domain, which helps fill the gap and enables new research into CTI summarization techniques.
3. Data Construction
As shown in Figure 2, the proposed multi-stage annotation pipeline consists of four stages. The first stage is collecting data from different sources, namely open source report (CyberMonitor and et al., 2019), threat encyclopedia (enc, 2020), and security company report (sym, 2020) to ensure the diversity and coverage. The second stage is parsing and cleaning which aims to parse original PDF reports into readable text, then obtain high-quality input data by manually cleaning and filtering. Next follows the prompt schema stage, where domain experts design various prompts based on different tasks, to fully utilize the powerful text generation capabilities of LLMs and prepare for the annotation. Finally, the intelligence summarization stage. To combine system productivity and expert experience efficiently, we transform the manual annotation process into ranking and classifying, having experts review and rank the outputs of LLMs to obtain the final summary. Next, we first introduce the multi-stage annotation pipeline, and then make a statistics and analysis of the CTISum.
3.1. The Multi-stage Annotation Pipeline
The multi-stage annotation pipeline focuses on obtaining the high-quality CTISum dataset, where we attempt to combine the strengths of both expert experience and LLMs (like GPT-4 (OpenAI, 2023) or Claude2 (Claude, 2023a)) to achieve semi-automatic summarization annotation for efficiency in the cybersecurity domain.
@@ 3.1. The Multi-stage Annotation Pipeline
To gather relevant CTI reports and build CTISum, we use the reports scraped from various sources, including open source report like APTnotes [^1] https://github.com/blackorbird/APT_REPORT, threat encyclopedia like Trendmicro [^2] https://www.trendmicro.com, and security company report like Symantec [^3] https://www.broadcom.com/support/security-center to ensure the diversity and coverage of the CTISum. The range of collected data is covered from 2016 to 2023. For the summarization tasks, a balanced dataset of 1,345 documents is sampled and annotated with two categories: CTIS and APS.
@@
The purpose of this stage is to develop a detailed prompt schema to guide the LLMs ( GPT-4, Claude2 and ChatGLM3) in producing candidate summaries, such as highlighting key events and summarizing reports, attack processes, some useful annotations, etc. To generate accurate and consistent summaries for the CTIS task and the APS task, CTI experts analyze the parsed and cleaned text and then design the CTI prompt schema and attack process prompt schema, respectively. As shown in Figure 2, several key annotation prompts are defined as follows:
@@
To combine system productivity and expert experience more efficiently, we transform the manual annotation process into ranking and classifying model-generated results, having experts review and rank the outputs into overall summaries. Specifically, LLMs like GPT-4 (OpenAI, 2023), Claude2 (Claude, 2023a), ChatGLM3 (Zeng et al., 2022) are leveraged to generate multiple coarse-grained summaries for the proposed two tasks. Then, domain experts are involved in reviewing and ranking the candidate summaries, selecting the better content to combine into a final gold summary. If the generated summaries turn out to be inadequate, the example would be discarded. The domain experts play a key role in evaluating and improving the model-generated summaries to produce a high-quality benchmark. Their human judgment and domain expertise complement the capabilities of LLMs. This collaborative human-AI approach allows the creation of abstractive summaries that are coherent, relevant, and accurate.
Table 1. Statistics of CTISum dataset and existing different domain summarization datasets. Numbers which are not reported are left blank. The numbers for the datasets marked with $\ast$ are copied from the paper (Mukherjee et al., 2022), whereas the ones marked with $\dagger$ are copied from the work (Narayan et al., 2018). Doc.: document, Avg.: average, Len.: length, APS: attack process summarization.
@@
Arxiv/PubMed (Cohan et al., 2018$\ast$ 346,187 5,179.22 257.44 - No Academic
CNN (Nallapati et al., 2016$\dagger$ 92,579 760.50 45.70 - No News
DailyMail (Nallapati et al., 2016$\dagger$ 219,506 653.33 54.65 - No News
XSum (Narayan et al., 2018$\dagger$ 226,711 431.07 23.26 - No News
BookSum Chapters (Mukherjee et al., 2022$\ast$ 12,630 5,101.88 505.42 - No Books
@@
GovReport (Huang et al., 2021) 19,466 9,409.40 553.40 - No Government
Discharge (Zhu et al., 2023) 50,000 2,162.29 28.84 - No Biomedicine
Echo (Zhu et al., 2023) 162,000 315.30 49.99 - No Biomedicine
ECTSum (Mukherjee et al., 2022) 2,425 2,916.44 49.23 - No Finance
IN-Ext (Shukla et al., 2022) 50 5,389 1,670 - No Court Rulings
UK-Abs (Shukla et al., 2022) 793 14,296 1,573 - No Court Rulings
@@
Table 2. Fleiss Kappa evaluation of CTISum dataset.
Score 1 2 3 4 5 Fleiss Kappa
Result 0 9 58 129 54 0.37
Score 1 and 2 3 4 and 5 Fleiss Kappa
Result 9 58 183 0.61
3.2. Data Statistics
Table 1 describes the statistics of the CTISum dataset and existing summarization datasets from different domains. The datasets are divided into general and specific domain categories, with the latter attracting substantial interest in recent years from 2021 to 2023. And we make the first attempt to build the CTI summarization dataset in the cybersecurity domain. Specifically, CTISum has 1,345 documents from different sources, with an average document length of 2,865 words, with an average CTI summary length of 200 words, and with a high compression ratio of 14.32. Notably, CTISum is the only dataset with the subtask, while other datasets simply make a general summary of the document. The average length of the attack process summaries is about 118 words, which is shorter than the general summaries. This highlights the challenging nature of generating focused summaries that capturing the attack process from long documents. In addition, the Fleiss Kappa evaluation of the dataset can be seen in Sec. 3.3.

### v2

@@ Abstract
Cyber Threat Intelligence (CTI) summarization involves generating concise and accurate highlights from web intelligence data, which is critical for providing decision-makers with actionable insights to swiftly detect and respond to cyber threats in the cybersecurity domain. Despite that, the development of efficient techniques for summarizing CTI reports, comprising facts, analytical insights, attack processes, and more, has been hindered by the lack of suitable datasets. To address this gap, we introduce CTISum, a new benchmark dataset designed for the CTI summarization task. Recognizing the significance of understanding attack processes, we also propose a novel fine-grained subtask: attack process summarization, which aims to help defenders assess risks, identify security gaps, and uncover vulnerabilities. Specifically, a multi-stage annotation pipeline is designed to collect and annotate CTI data from diverse web sources, alongside a comprehensive benchmarking of CTISum using both extractive, abstractive and LLMs-based summarization methods. Experimental results reveal that current state-of-the-art models face significant challenges when applied to CTISum, highlighting that automatic summarization of CTI reports remains an open research problem. The code and example dataset can be made publicly available at https://github.com/pengwei-iie/CTISum.
keywords:
[…]
{highlights}
To the best of our knowledge, we make the first attempt to build a new benchmark CTISum with the CTIS task and a novel APS subtask in the cybersecurity domain.
A multi-stage annotation pipeline is designed to obtain the high-quality dataset with the assistance of LLMs while manually controlling the quality.
Experiments on CTISum demonstrate the challenge of the proposed two tasks, meanwhile indicating a large space for future research.
1 Introduction
@@
Cyber Threat Intelligence (CTI), also known as threat intelligence, is knowledge, skills and experience-based information concerning the occurrence and assessment of both cyber and physical threats as well as threat actors [17, 12]. The CTI data often originates from diverse web sources, including forums, blogs, and open-source repositories, making it difficult for analysts to efficiently locate the most relevant and high-value intelligence from the web. Therefore, it becomes crucial to automatically summarize the knowledge contained in CTI reports, which could help analysts simply identify events, patterns, cyber attacks and conclusions. Furthermore, CTI summarization has broad applicability across domains like cybersecurity [36], military intelligence [39], technical alerts, threat modeling, and more, where complicated information needs to be distilled into concise insights.
@@
Figure 1: An example in the CTISum. Blue arrow and orange arrow indicate the CTIS and APS tasks, respectively. Red font represents the key words.
@@
With the ongoing advancement of summarization techniques, numerous datasets have emerged, such as CNN/DailyMail [28] and XSum [29], among others. In addition, the introduction of these datasets further drives the development of associated technologies like PGNet [37], BART [19], BRIO [23], etc. Furthermore, domain-specific summarization researches have gradually arisen. Examples include biomedicine [46], finance [26], law [38], etc. This further facilitates various applications involving aiding medical decision-making, generating financial reports, and summarizing legal documents. The boom of domain-specific summarization promotes the understanding of domain terminology and context, which is crucial for generating high-quality summaries for further research.
@@
Despite remarkable progress made in automatic text summarization, effective methods for summarizing CTI reports in the cybersecurity domain remain largely unexplored. This is primarily due to the lack of available dataset. The unique characteristics of CTI reports, such as technical jargon, evolving threat landscapes, longer reports and fragmented data information, make it challenging to create an annotation, which in turn hinders the development of robust models and evaluation benchmarks. However, recent advances in Large Language Models (LLMs) [14, 9, 30, 40, 7, 6] have shown promising text understanding and generation capabilities. Therefore, this raises the question: How might we construct a high-quality CTI summarization dataset with the assistance of LLMs?
@@
In this paper, we construct a new benchmark CTISum based on threat intelligence obtained from diverse web sources in the cybersecurity domain. In addition to the CTI Summarization (CTIS) task, considering the importance of attack process, we propose a novel fine-grained subtask, Attack Process Summarization (APS), to enable defenders to quickly understand reported attack behaviors, assess risks and identify security vulnerabilities. Specifically, a multi-stage annotation pipeline is designed to preprocess web CTI reports and obtain annotation data with the assistance of LLMs, which includes the data collection stage, parsing & cleaning stage, prompt schema stage and intelligence summarization stage. To further keep the high quality of the CTISum, we additionally employ expert reviews for double checking (details can be seen in Section 3.1). With proper data, automated systems can help synthesize lengthy CTI reports into concise and accurate summaries for different intelligence analysts. A sampled report and reference summary in CTISum are shown in Figure 1. The system not only requires to make a general summary of the full document (task 1), but also needs the capability to focus on and generate the attack process (task 2). After cleaning and human checking, CTISum obtains 1,345 documents and corresponding summaries.
@@
What makes CTISum a challenging dataset can be presented as follows. First, the average document length is about 2,865 words. Current deep learning techniques are incapable of processing documents surpassing 512/1024 tokens in length (after tokenization), such as GPT1[33], T5[34], BART[19] and so on. Although LLMs can deal with longer input, however, fine-tuning LLMs is resource consuming. Moreover, LLMs are generally pre-trained on generic datasets, so they often fall short when performing zero-shot tasks in specific domains, such as cybersecurity. Second, the document-to-summary compression ratio on the two tasks are 14.32 (2865.60 / 200.04) and 22.23 (2865.60/118.27), respectively (details in Table 1). The high compression ratio shows competitiveness with the current long document summarization datasets, requiring systems to be extremely precise in capturing only the most relevant facts from lengthy documents in a minimal number of words. Finally, CTISum has another subtask, APS, which means the system should be capable of the ability to capture the fine-grained attack process described in CTI reports.
@@
Figure 2: The overview of the proposed multi-stage annotation pipeline for CTIS and APS tasks.
2 Related Work
2.1 Text Summarization
A considerable amount of existing researches in automatic text summarization [35, 28, 27, 4, 18, 44] have explored to improve summary generation for news articles, using popular datasets such as CNN/DailyMail [28], XSum [29], etc. In addition, scientific document summarization has emerged as another critical area, with datasets based on ArXiv/PubMed [8], and other academic sources. The main approaches cover two aspects. First, extractive methods [13, 5, 45, 21, 22, 44] involve identifying and extracting key sentences or passages from the original text to form the summary. For instance, Liu et al. [22] introduce a document-level encoder based on BERT and propose a general framework for summarization. Second, abstractive techniques focus on generating new sentences to capture the meaning of the source text. For example, sequence-to-sequence models with attention [25, 27, 31] are usually utilized to generate the probability distributions on the vocabulary. In addition, See et al. [37] present a hybrid pointer generator architecture for abstractive summarization. And more recently, transformer-based models like T5 [34], BART [19] and Longformer[3] have shown advanced performance on the summarization task.
@@ 2.1 Text Summarization
2.2 Domain-specific Summarization
For the domain-specific summarization, there are many potential applications across different domains like government [16], finance [26], court rulings [38] and biomedicine [46]. For example, in the government domain, summarization systems can help process large volumes of regulatory text and legal documents. Huang et al. [16] propose a novel efficient encoder-decoder attention with head-wise positional strides to effectively double the processed input sequence size. For the finance industry [26], summarization of earnings reports, financial statements, and news can assist investors and analysts in making decisions. In addition, summarizing court rulings, case law, and litigation documents [38] helps legal professionals in their work. Within biomedicine [46], healthcare professionals can make informed decisions by using summaries of clinical trial findings, medical literature, and patient records. Different from the above domains, we make the first attempt to build the CTI summarization dataset in the cybersecurity domain, which helps fill the gap and enables new research into CTI summarization techniques.
3 Data Construction
As shown in Figure 2, the proposed multi-stage annotation pipeline consists of four stages. The first stage is collecting data from diverse web sources, namely open source report [10], threat encyclopedia [2], and security company report [1] to ensure the diversity and coverage. The second stage is parsing and cleaning which aims to parse original PDF reports into readable text, then obtain high-quality input data by manually cleaning and filtering. Next follows the prompt schema stage, where domain experts design various prompts based on different tasks, to fully utilize the powerful text generation capabilities of LLMs and prepare for the annotation. Finally, the intelligence summarization stage. To combine system productivity and expert experience efficiently, we transform the manual annotation process into ranking and classifying, having experts review and rank the outputs of LLMs to obtain the final summary. Next, we first introduce the multi-stage annotation pipeline, and then make a statistics and analysis of the CTISum.
3.1 The Multi-stage Annotation Pipeline
The multi-stage annotation pipeline focuses on obtaining the high-quality CTISum dataset, where we attempt to combine the strengths of both expert experience and LLMs (like GPT-4 [30] or Claude2 or 3.5 [6]) to achieve semi-automatic summarization annotation for efficiency in the cybersecurity domain.
@@ 3.1 The Multi-stage Annotation Pipeline
To gather relevant CTI reports and build CTISum, we use the reports scraped from diverse web sources, including web open source report like APTnotes [^1] https://github.com/blackorbird/APT_REPORT, threat encyclopedia like Trendmicro [^2] https://www.trendmicro.com, and security company report like Symantec [^3] https://www.broadcom.com/support/security-center to ensure the diversity and coverage of the CTISum. The range of collected data is covered from 2016 to 2024. For the summarization tasks, a balanced dataset of 1,345 documents is sampled and annotated with two categories: CTIS and APS.
@@
The purpose of this stage is to develop a detailed prompt schema to guide the LLMs (Claude2 or 3.5, ChatGLM3, etc.) in producing candidate summaries, such as highlighting key events and summarizing reports, attack processes, some useful annotations, etc. To generate accurate and consistent summaries for the CTIS task and the APS task, CTI experts analyze the parsed and cleaned text and then design the CTI prompt schema and attack process prompt schema, respectively. As shown in Figure 2, several key annotation prompts are defined as follows:
@@
To combine system productivity and expert experience more efficiently, we transform the manual annotation process into ranking and classifying model-generated results, having three experts review and rank the outputs into overall summaries. Specifically, LLMs like GPT-4o [30], Claude2 or 3.5 [6], ChatGLM3 [42] are leveraged to generate multiple coarse-grained summaries for the proposed two tasks. Then, three domain experts are involved in reviewing and ranking the candidate summaries, discussing and selecting the better content to combine into a final gold summary. If the generated summaries turn out to be inadequate, the example would be discarded. The domain experts play a key role in evaluating and improving the model-generated summaries to produce a high-quality benchmark. Their human judgment and domain expertise complement the capabilities of LLMs. This collaborative human-AI approach allows the creation of abstractive summaries that are coherent, relevant, and accurate.
Table 1: Statistics of CTISum dataset and existing different domain summarization datasets. Numbers which are not reported are left blank. The numbers for the datasets marked with $\ast$ are copied from the paper [26], whereas the ones marked with $\dagger$ are copied from the work [29]. Doc.: document, Avg.: average, Len.: length, APS: attack process summarization.
@@
Arxiv/PubMed [8$\ast$ 346,187 5,179.22 257.44 - No Academic
CNN [28$\dagger$ 92,579 760.50 45.70 - No News
DailyMail [28$\dagger$ 219,506 653.33 54.65 - No News
XSum [29$\dagger$ 226,711 431.07 23.26 - No News
BookSum Chapters [26$\ast$ 12,630 5,101.88 505.42 - No Books
@@
GovReport [16] 19,466 9,409.40 553.40 - No Government
Discharge [46] 50,000 2,162.29 28.84 - No Biomedicine
Echo [46] 162,000 315.30 49.99 - No Biomedicine
ECTSum [26] 2,425 2,916.44 49.23 - No Finance
IN-Ext [38] 50 5,389 1,670 - No Court Rulings
UK-Abs [38] 793 14,296 1,573 - No Court Rulings
@@
3.2 Data Statistics
Table 1 describes the statistics of the CTISum dataset and existing summarization datasets from different domains. The datasets are divided into general and specific domain categories, with the latter attracting substantial interest in recent years from 2021 to 2024. And we make the first attempt to build the CTI summarization dataset in the cybersecurity domain. Specifically, CTISum has 1,345 documents from diverse web sources, with an average document length of 2,865 words, with an average CTI summary length of 200 words, and with a high compression ratio of 14.32. Notably, CTISum is the only dataset with the subtask, while other datasets simply make a general summary of the document. The average length of the attack process summaries is about 118 words, which is shorter than the general summaries. This highlights the challenging nature of generating focused summaries that capturing the attack process from long documents. In addition, the Fleiss Kappa evaluation of the dataset can be seen in Sec. 5.5.